OpenStudy (anonymous):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

OpenStudy (anonymous):

this is why mathemeticians invented fraction notation do you mean \[\frac{3}{5}n-\frac{4}{5}=\frac{1}{5}n\] like that?

OpenStudy (anonymous):

lol yes sorry

OpenStudy (anonymous):

no problem fractions are annoying get rid of them by multiplying everything by 5, i.e. rewrite is with no denominators

OpenStudy (anonymous):

clear, or no?

OpenStudy (anonymous):

yes is the answer -1?

OpenStudy (anonymous):

or did i do it wrong

OpenStudy (anonymous):

i don't think so solve \[3n-4=n\]

OpenStudy (anonymous):

i mean "i don't think the answer is -1"

OpenStudy (anonymous):

would it be one of the fractions

OpenStudy (anonymous):

just do it, as they say

OpenStudy (anonymous):

i.e. don't guess, solve the equation

OpenStudy (anonymous):

i got it its 2